What is the shortest distance in which an automobile travelling on a 100km/hr can be stopped. (a) when the friction between the tyre and the road is 0.5 (b) on a wet day when the coefficient of friction is only 0.3

1
Expert's answer
2021-11-29T11:39:00-0500

"s=\frac{v^2}{2a}"


"F=ma\to a=F/m=\mu mg/m=\mu g"


"100\ km/hr=27.8\ m/s"


(a) "s=\frac{v^2}{2\mu g}=\frac{27.8^2}{2\cdot 0.5\cdot9.8}=78.7\ (m)"


(b) "s=\frac{v^2}{2\mu g}=\frac{27.8^2}{2\cdot 0.3\cdot9.8}=131.4\ (m)"